OpenStudy (anonymous):

Question Part Points Submissions Used Find an equation of the line that passes through the point and has the indicated slope m. (Let x be the independent variable and y be the dependent variable.) (5, −3); m = −1

OpenStudy (anonymous):

for this question you will use the point slope formula since you have a point and a slope. the formula is (Y-y1)= m(X-x1)...so it is y+3=-1(x-5). you just then simplify this by distributive property.